OpenStudy (anonymous):

For the Integral (1/X^p) dx from 0 to 1, find the values of p for which the integral converges ?

OpenStudy (anonymous):

are you sure about the boundaries ?

OpenStudy (anonymous):

p not 1

OpenStudy (turingtest):

i mean p not equal to 1

OpenStudy (anonymous):

p is not =1

OpenStudy (anonymous):

Yes :(

OpenStudy (turingtest):

Why the frown? That's the answer. \[p \neq1\]

OpenStudy (anonymous):

it says find the value of p ?

OpenStudy (anonymous):

values of p .. so for p!= 1 converge

OpenStudy (anonymous):

how do u prove it

OpenStudy (turingtest):

the question wants all values of p that don't converge for that function \[\int\limits\limits_{0}^{1}1/x^pd=x^(1-p)/(1-p)\] evaluated from 0 to 1 so as long as p isn't 1 you don't get a zero on the denominator, hence the integral is finite (converges)